OpenStudy (vortish):

by what number must both sides be multiplied in order to solve 5/7y =8

OpenStudy (anonymous):

solve for y?

OpenStudy (vortish):

its says what both sides be multiplied in order to solve the equation 5/7y = 8

OpenStudy (anonymous):

can you clarify if it is "five sevenths times y equals 8" ? Or is it 5/(7y) = 8 ? In other words, is the y in the denominator?

OpenStudy (vortish):

no its 5/7 y is for both = 8\[\frac{ 5 }{ 8 }y=8\]

OpenStudy (vortish):

its seven not 8

OpenStudy (anonymous):

so you need to multiply both sides by a number that will cause y to be all by itself on the left after... they are wanting you to think about inverses of fractions... if you have (5/7) y on the left, you can multiply by the inverse of (5/7) which is (7/5) So (5/7) y = 8 Multiply both sides by the inverse of the fraction in front of the "y"... (5/7)(7/5) y = (8)(7/5) then the fractions on the left cancel out, leaving just y y = 8(7/5) = 56/5

OpenStudy (vortish):

ok that makes sense this is a linear equation so it makes sense

OpenStudy (anonymous):

ok, glad to help :) Even if the y-term didn't look like it had a fraction, like 8y, you could still treat that like a fraction like (8/1)y... then you would multiply both sides by the inverse, which would be (1/8) to cancel the stuff in front of the "y". That's obviously the same thing, in different words, to saying you divide both sides of the equation by 8 to solve for y if you started with 8y.

OpenStudy (vortish):

you rock